OpenStudy (anonymous):

Three balls are drawn from a box containing 6 red marbles, 4 white marbles and 5 blue marbles. Find the probability that they are drawn in the order: red, white, and blue if each ball is drawn with replacement

OpenStudy (anonymous):

are you currently covering conditional probability?

OpenStudy (anonymous):

yep u got preparing for exams

OpenStudy (anonymous):

I'm not sure if it is P(red AND white AND blue) or if it is P(red|white|blue)

OpenStudy (anonymous):

oh, one sec i think i got this.

OpenStudy (anonymous):

P(pulling a red on the first one) = 6/15

OpenStudy (zehanz):

Draw them one by one, each time putting it back. You want to know: P(r,w,b). So: Draw 1st. P(red)=6/15. Draw 2nd. P(white)=4/15. Draw 3rd. P(blue)=5/15. Multiply those probabilities to find P(r,w,b).